OpenStudy (anonymous):

how can you write the expression with a rationalized denominator? ^3sqrt2/^3sqrt4 ??

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle \frac{\sqrt[3]{2} }{\sqrt[3]{4}} }\) like this?

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

Example problem 1: \(\large\color{#000000 }{ \displaystyle \frac{\sqrt[3]{3} }{\sqrt[3]{9}} }\) I know that, \(\large\color{#000000 }{ \displaystyle \sqrt[3]{9} \times \sqrt[3]{3}=\sqrt[3]{27}=3 }\) So, let's multiply times \(\large\color{#000000 }{ \displaystyle \sqrt[3]{3}}\) on top and bottom. \(\large\color{#000000 }{ \displaystyle \frac{\sqrt[3]{3} \times \sqrt[3]{3} }{\sqrt[3]{9}\times \sqrt[3]{3}} =\frac{\sqrt[3]{9}}{3} }\)

OpenStudy (anonymous):

Ok

OpenStudy (solomonzelman):

Example problem 2: \(\color{#000000 }{ \displaystyle \frac{\sqrt[3]{4} }{\sqrt[3]{16}} }\) I know that, \(\color{#000000 }{ \displaystyle \sqrt[3]{16} \times \sqrt[3]{4}= \sqrt[3]{4^3}=4 }\) So, let's multiply times \(\large\color{#000000 }{ \displaystyle \sqrt[3]{4}}\) on top and bottom. \(\color{#000000 }{ \displaystyle \frac{\sqrt[3]{4} \times \sqrt[3]{4} }{\sqrt[3]{16}\times \sqrt[3]{4}} =\frac{\sqrt[3]{16}}{4} }\)

OpenStudy (solomonzelman):

you are just multiplying (on top and bottom) times something that would make the denominator into integer, basically/

OpenStudy (anonymous):

Ooh ok. So you would multiply the top by the bottom and get ^3sqrt 8

OpenStudy (solomonzelman):

Wow

OpenStudy (anonymous):

Wouldn't it just be 8?

OpenStudy (solomonzelman):

Well, you are multiplying by \(\sqrt[3]{2}\), but not because it is the top, RATHER you are doing because that would (indeed) get you \(\sqrt[3]{8}=2\). And, you are multiplying times \(\sqrt[3]{2}\) on TOP and BOTTOM.

OpenStudy (anonymous):

Oh ok.

OpenStudy (solomonzelman):